OpenStudy (anonymous):

(-5,2) (3,-4) find slope intercept form

OpenStudy (anonymous):

Hint: use two point fotmula to calculate the equation of the line

OpenStudy (anonymous):

i did and i got y= -3/4-1.75

OpenStudy (anonymous):

y-2=(-4-2)/(3+5) (x+5) 8y-16=-6x-30 8y=-6x-14 y=-3/4 x - 7/4

OpenStudy (anonymous):

Thats the correct answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome